A car can be purchased on credit by payments of 410 euro each month, over a period of 36 months. Calculate the price of the car if purchased on credit. The cash price for the car is 14,000. Calculate the extra cost of the credit as a percentage of the cash price, correct to 2 decimal places.

An i-pod can be purchased for 380 dollars cash. Alternatively, it can be paid for over a year by monthly instalments of 34 dollars. Calculate the extra cost of the credit option. Express this cost as a percentage of the cash price, correct to 1 decimal place.

Answer by rfer(16322)

410*36=14760
760/14000=.05=5%
--------
34*12=408
28/380=.07=7%=